OpenStudy (anonymous):

Evaluate the integral by using the following transformation: ∫∫(R) x*y^2 dA, where R is the region bounded by the lines x-y=2, x-y=-1, 2x+3y=1, and 2x+3y=0; let x=1/5(3u+v), y=1/5(v-2u)

OpenStudy (turingtest):

so, where are you stuck?

OpenStudy (anonymous):

Well, I tried doing the problem and I didn't get the right answer. Are these the correct steps to get to the answer? 1. Plot x/y coordinate graph. 2. Convert critical xy points to uv points. 3. find the Jacobian. 4. Set up the integral in terms of u and v with the jacobian.

OpenStudy (anonymous):

I do have the right answer if you want it.

OpenStudy (turingtest):

your steps are fine, but I would just convert the lines to u and v instead of the "critical xy points"

OpenStudy (anonymous):

oh? How would I do that? That sounds a lot easier.

OpenStudy (turingtest):

just sub in x=1/5(3u+v), y=1/5(v-2u) into each equation for the boundary of R

OpenStudy (anonymous):

Ohhhh, alright. So once I converted to uv form would I just replot the graphs and look for my bounds?

OpenStudy (turingtest):

yeah, it will most likely be a square in the u v plane

OpenStudy (anonymous):

Gotcha, Ill rework the problem. That sounds a lot better than just looking for the points which takes forever.

OpenStudy (turingtest):

good luck!

OpenStudy (anonymous):

thanks